OpenStudy (vannah):

what is the slope of the line that passes through the points (-5,4)and (15,-4)??

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (-4-4)/(15-(-5)) m = (-8)/(20) m = -2/5

OpenStudy (vannah):

how you ghott 20

jimthompson5910 (jim_thompson5910):

15-(-5) = 15+5 = 20 Remember, subtracting a negative is like adding. Eg: I owe you $5. So I have 5 IOUs. If I subtract 5 IOUs, then I'm essentially adding $5 to my pocket.

OpenStudy (vannah):

ooo okayy

OpenStudy (vannah):

i give you a medall for that and thankss again